OpenStudy (anonymous):

Find the inverse of each function. Is the inverse a function? Please help me with this problem?

OpenStudy (anonymous):

@jim_thompson5910 can you explain this to me?

jimthompson5910 (jim_thompson5910):

To find the inverse, swap x and y, then solve for y \[\Large y = x^2 + 2\] \[\Large x = y^2 + 2\] \[\Large x-2 = y^2\] \[\Large y^2 = x-2\] \[\Large y = \pm\sqrt{x-2}\] \[\Large f^{-1}(x) = \pm\sqrt{x-2}\] Because y is effectively two functions at the same time (plus/minus of \(\Large \sqrt{x-2}\)), there is no way y can be a single function. Also, graphing this function will show us that it fails the horizontal line test. So the inverse is NOT a function.